Subject: 3.5-m Users Committee minutes 5/14/07

From: Bruce Gillespie

Submitted: Thu, 17 May 2007 15:26:41 -0600

Message number: 1028 (previous: 1027, next: 1029 up: Index)

                         APO 3.5-m Users Committee Phonecon, 5/14/07

Attending:  Suzanne Hawley, Bruce Gillespie, Mark Klaene, Remy  
Indebetouw, Russet McMillan, Michael Strauss, Jon Holtzman, John Bally

Absent: Scott Anderson, Jon Fulbright, Al Harper

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton (Strauss) - Michael said everything is going fine with the  
PU users.

JHU (Fulbright) - no report.

Chicago (Harper) - no report.

Washington (Suzanne) - Suzanne reported that a run with Agile is  
about to be underway, which will take photometric data jointly with  
other observatories.

NMSU (Holtzman) - Jon had nothing to report.

Virginia (Indebetouw) - nothing to report, except that UVa users have  
been getting data recently in spite of the weather.

Colorado (Bally) - John mentioned that John Stock has been getting  
data, and that he (Bally) is collaborating on a GFP run that is  
currently underway.


Upcoming NIC-FPS availability, plans - Bally:

John reported that the plan is to take NIC-FPS to Boulder on 24 July  
to effect some upgrades and modifications.  This will have impact to  
Q3 scheduling of the instrument on the telescope, because the  
instrument will not return to APO until September, nominally.   
Schedulers have been already been told this, and users should also  
note this when making their Q3 requests.

Work planned includes modifying the detectors (both science and  
engineering) for 16-channel readouts.  In parallel, the NIR Fabry- 
Perot will be tested and commissioned in the lab.  The set point for  
the operational temperature of the etalon is an open question, and if  
modifications are necessary to operate it at a warmer temperature are  
required, the return schedule to APO may be at some risk.


Discussion of telescope/instruments report - Klaene:

The report is given below; Mark walked through the highlights, and  
there was little discussion:


     3.5-m Telescope, Instruments, and CIF Projects Highlights,  
4/14/07 through 5/10/07

                                  Mark Klaene

0) Overview

Spring weather continues to be intermixed with mild winter storms.   
Recent telescope closures have been primarily due to heavy cloud  
cover and precipitation, atypical for this time of year.  Fire danger  
is unlikely to reach extreme levels this spring, and restrictions and  
forest closures are not likely as we continue with moderate rain and  
high humidity along with cooler temperatures.  Fir trees near the  
observatory are being attacked by defoliator insects.  We are working  
with the USFS on determining the effects and possible actions if  
these bugs reach the observatory grounds.

1) Telescope

No significant problems this period. Subtle az and altitude tracking  
variations are being monitored; they are so far  having little to no  
effect on operations.

2) Instruments

   o Agile is preparing for its second run.  The "actual" (i.e., not  
the substitute) Agile camera is back after repair and ready for the  
run. Minor mounting and handling improvements have been made.  The  
fiber-optic interface is still not working and will require further  
bench testing back at UW.  The focal reducer is not yet available.

   o APOLLO interlock problems have been corrected.

   o The modified grating holder for the DIS gratings was tested and  
found to be less repeatable than desired. A modification is now in work.

   o The spare Leach controller for NIC-FPS and TripleSpec has been  
ordered.  Planning to remove NIC-FPS from mid-July through mid-Sept  
for upgrades is being finalized.

   o DIS red camera had computer interface card failure in its ICC.   
A spare was on hand and the card was replaced. No sky time was lost.  
This is a commercial card and will be repaired for future use as a  

   o Assimilation of SPIcam into TUI is planned to start in mid-May.   
The issue of SPIcam TUI scripts was resolved by having Bill Ketzeback  
talk to Russell Owen.  When SPIcam TUI interface will be ready for  
user testing is TBD.

3) Engineering and CIF projects

   o The motor and interface electronics for the NA2 rotator upgrade  
are being purchased.  A preliminary motor mount design is in review.

   o Most of the parts are on hand for phase one of the upgrades to  
the axis controllers.  Much of the software is completed and running  
in a simulation mode.  Plan is to integrate the software with the NA2  
rotator this summer, and pending successful operation, begin  
integration with the drive axes in late summer/fall, with the project  
completed by the end of the year.

4) Miscellaneous

   o Truss lamp replacement project is still in the research phase.

   o Work on TripleSpec purge air and heat exhaust is complete until  
instrument arrival.

   o New GPS clocks are on site and running, final certification  
testing is in work.

   o Summer shutdown still planned for Aug 20-Sept 16.


Other misc. topics - Group:

   o Suzanne reiterated that the SPIcam/TUI upgrade is starting this  
week. and the SPIcam TUI interface will likely be ready for user  
testing in the next few weeks.  SPIcam users with time in June should  
be prepared to use the new TUI interface, and to provide feedback on  
its operation.

   o Russet said that she was able to get some drop-in engineering  
observing tests recently, and she got some FastCam baseline data that  
can be compared to similar observations that will be taken after the  
drive boxes are swapped out with spares at the end of the month.   
Also, the az and rotator drives have been over-currenting recently;  
Mark thinks this is currently a "feature" of the system related to  
slow tracking speeds at certain pointing angles, and will need to  
wait until July for serious attention unless the problem gets worse  
in the meantime.

   o Suzanne reported that there is a newly formed NSF "small"  
telescope committee (i.e., less than 6.5-meters aperture) that is  
looking at science and coordination issues related to the use and  
refurbishment of these telescopes.  This group is responding to the  
Senior Review recommendation for providing more access to the  
community to small and mid-sized telescopes.  They are meeting  
several times this year, and will be soliciting community input. This  
could ultimately lead to an NSF program similar to the TSIP program,  
but for the smaller telescopes. There is no immediate effect for us  
but Suzanne feels that it could be a source of future resources for  
our telescope if our budget becomes tight. She is on the committee as  
a liaison to ACCORD (the AURA large telescope directors committee)  
and will distribute information about providing input when it becomes  

   o Suzanne reminded us of our student training policy.  Students  
are required to be accompanied at the site by their advisors for the  
three day training and orientation visit; see 
35m_operations/35m_manual/General_Information/Training.html.  She  
will consider exceptions to this rule on a case-by-case basis, with  
compelling reasons.  We may also consider revising this policy when  
and if it makes sense to do so.


Actions from earlier meetings - group

[open from previous months]:


[new actions from last month]:

ACTION: Suzanne will follow up with Russell Owen about TUI  
scripting.  Users are encouraged to provide feedback to their users  
reps about integrating Spicam into TUI.
STATUS: Russell is willing to include some modest number of useful  
TUI scripts when SPIcam is implemented into TUI.  No feedback has  
been received, so we plan to close this action item.

DISCUSSION:  Jon Holtzman will contact Russell to discuss the details  
of TUI scripts for SPIcam.

UPDATE:  Action item reopened (was closed last month). Not sure this  
has been adequately communicated to the users; see item in telescope/ 
instrument reports above.  Jon Holtzman agreed at the telecon to talk  
to Bill Ketzeback and Russell Owen about this.  This action is now re- 


Next meeting:

The next Users Committee phonecon will be on Monday, 11 June, at 8:30  
AM Pacific Time.  Agenda and other materials will be sent to the  
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1028 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to